R=0.082 L atm mol^-1 K^-1

The pressure in atm, if 8.25 X 10^-2 mole gas occupies a volume of 255 mL at 115ºC.
1
Expert's answer
2020-06-25T07:52:53-0400

pV=nRT;

p=nRT/V=8.25*10^(-2) * 8.314 * 388/0.255 =1043.65 Pa= 0.01 atm

Answer: 0.01 atm